### v1

# GPU Cluster Scheduling for Network-Sensitive Deep Learning

Abstract.
Deep learning (DL) constitutes a significant workload within public or private datacenters. Organizations commonly engage in training multiple Deep Neural Networks (DNNs) in multi-tenant GPU clusters. The primary consideration in training such models is the associated ”cost”, which correlates directly with the GPU usage duration in the cluster. Previous studies have demonstrated that communication overheads of interconnected GPU clusters could be a significant component of the overall training time and, thus, minimizing the communication overhead with intelligent scheduling of training jobs on physically close GPUs should minimize the long training times. However, state-of-the-art cluster schedulers are rather agnostic to the proximity-based job consolidation. To compound the issue, different DNN models display varied communication overheads even with similar levels of consolidation. This results in prolonged training times.
In this context, we propose a novel GPU-cluster scheduler for distributed DL (DDL) workloads that enables proximity-based consolidation of GPU resources based on the DDL jobs’ sensitivities to the anticipated communication-network delays. Our scheduler consists of three major components: (i) a classical delay scheduling algorithm to facilitate job placement and consolidation; (ii) a network-sensitive job preemption strategy; and (iii) an ”auto-tuner” mechanism to optimize delay timers for effective delay scheduling. Additionally, to enable a cost-effective methodology for large-scale experiments, we develop a data-driven DDL cluster simulation platform. Employing the simulation platform we compare against several state-of-the-art alternatives on real-world workload traces to demonstrate the benefits of our design. Our scheduler can provide improvement of up to 69% in end-to-end Makespan for training all jobs compared to the prevailing consolidation-based scheduling methods, while reducing the average job completion time by up to 83% and minimizing the communication overheads by up to 98% under congested networking conditions.
copyright: none
1. INTRODUCTION
Deep learning (DL) with large-scale cloud computing resources has facilitated the development of deep (very large) neural network (DNN) models, driving cutting-edge advancements in AI. These models, powering breakthroughs in domains such as autonomous vehicles (self-driving, ), multimedia processing (multimedia-DL, ) and generative AI (GAI, ), often require training hundreds of millions, if not billions or trillions, of parameters to achieve the desired accuracy. To expedite model convergence within a reasonable time-frame, parallel training paradigms based on data parallelism (pytorch_dist, ) and model parallelism (megatron, ) are employed, simultaneously utilizing multiple GPUs (or related ”neural processing” accelerators like TPUs (tpu, )) to train a single model. Furthermore, AI enterprises or groups may train hundreds or even thousands of such DNN models (philly_trace_analysis, ). In certain cases, a Machine Learning as a Service (MLaaS) (mlaas, ) provider could aggregate a substantial number of DL jobs from its diverse clientele into a single extensive GPU cluster. However, GPUs are costly, and the prolonged training times of these models lead to exorbitant training expenses.
To alleviate the financial burden of DL, DNN workloads are executed on ”shared” GPU clusters, such as those offered by public cloud providers or within enterprise data centers (or private clouds). In the case of public cloud services, the cost is determined by the wallclock time spent utilizing cloud VMs, while for shared data centers, the cost is proportional to the time spent using GPU resources.
Recent studies (pipedream, ; alibaba_pai, ; stash, ) have shown that communication overhead is a major contributor to prolonged training times (and increased costs) in Distributed DL (DDL). This stems from the fact that GPUs within a cluster are interconnected through various types of networks. During DNN training, collective communication algorithms such as all-reduce, scatter-gather, and others are employed to exchange tensors among GPU workers. This communication latency can vary significantly depending on the underlying network type connecting the GPUs. Generally, the more proximal (physically close) the GPUs of a job are to each other within the network, i.e., the more ”consolidated” they are, the lower the communication overhead.
A DDL job can be consolidated in three typical ways: (i) placing all worker GPUs on the same physical machine; (ii) locating all worker GPUs on the same rack interconnected via Infiniband (Infiniband, ) technology; or (iii) spreading the worker GPUs across the network through inter-rack connections [^1] Hereafter referred to as the ”network.”. Each of these ”network tiers” introduces significantly different communication delays, with the least overhead on the same machine, followed by the rack, and then the network. To mitigate communication overhead, a straightforward approach involves consolidating all GPUs of a job as much as possible. However, due to multi-tenancy, when cluster managers attempt to consolidate a job on physically proximal GPUs, they often encounter delays while waiting for resources to become available, leading to increased queueing times. Furthermore, previous research (tiresias, ; stash, ) has demonstrated that DNN models exhibit varying degrees of ”network sensitivity,” implying that each DNN model would experience a different amount of increase in communication overhead as network conditions deteriorate. Thus, not all models would equally benefit from communication overhead reduction through consolidation but they would all incur the same queueing delay. Hence, it is crucial to consolidate DNN jobs based on their ”network sensitivities” i.e. relative communication overheads, in order to minimize queueing delays due to consolidation.
Prior cluster scheduler Tiresias (tiresias, ) implements a consolidation strategy for DDL jobs based on the model’s ”skew.” In this strategy, jobs with high skew are consolidated into as few machines as possible, while others are left unconsolidated. This approach may lack flexibility for ”moderate consolidation,” such as placing a job on the rack for a ”moderately network-tier” sensitive model. Moderate consolidation is now more viable for communication, thanks to the presence of contemporary networking hardware that offers high quality-of-service (QoS).
The contemporary Infiniband based (Infiniband, ) rack switch offers remarkably high bandwidth, reaching up to 400 Gb/s per port with the utilization of GPU RDMA (gpudirect-RDMA, ), effectively reducing rack-level communication overhead. Moreover, even modern network (Ethernet) based switches provide 800 Gb/s bandwidth per port, albeit with higher latency (nvidia-spectrum, ). Yet, DDL cluster schedulers remain oblivious to such advancements and are unable to harness the consequent potential performance improvements. To address these limitations, schedulers should be equipped with the ability to dynamically adjust consolidation based on the network sensitivity of individual jobs. This adjustment should take into account the diverse performance characteristics of different network tiers, leveraging the capabilities of modern networking hardware essential for lowering communication overhead.
To achieve these goals, we propose Dally, a holistic cluster scheduler for DL job placement that accounts for the underlying system network topology, as well as the network sensitivity of each job to make intelligent placement decisions. Dally accomplishes this by employing the classical delay scheduling algorithm (delay_scheduling, ) in the context of DDL workloads, coupled with a novel network-sensitive pre-emption strategy for job placement and consolidation. Specifically, for every DDL job, Dally consults its dedicated ‘delay timers’ to decide whether to accept a resource offer at the offered network-tier or wait in anticipation of an offer with better network-tier until a delay timer has elapsed. Traditionally, the delay timers are hand-tuned and remain fixed for a given system. However, in this work, we introduce a novel ‘auto-tuner’ that dynamically evolves with the system to tune these delay timers for high-efficacy delay scheduling. With the help of these innovations, Dally is able to outperform state-of-the-art schedulers in terms of makespan, job completion time (JCT) and communication overhead.
For evaluation of Dally, we have developed a simulation platform – ArtISt-sim. The use of simulations is motivated by the fact that development of novel DL cluster schedulers is currently impeded by the exorbitant costs associated with the large-scale experimental setups needed for such endeavors. A single GPU for DL costs thousands of dollars, and replicating a real-world DL data center necessitates a substantial number of these GPUs for extended periods.
ArtISt-sim has been developed by extending prior work (themis-nsdi, ) and utilizing ASTRA-sim (rashidi2020astra, ), a highly accurate DDL job simulator. Our simulation platform possesses two key capabilities: (i) it can dynamically determine network slowdowns based on the specific placement of a DDL job (its assigned GPUs) within the simulated cluster, and (ii) it can simulate a datacenter equipped with modern networking hardware (e.g., NVIDIA Quantum (nvidia-quantum, ) and Spectrum switches (nvidia-spectrum, )) which is crucial for this work. This combination of features enables ArtISt-sim to provide a realistic and accurate representation of modern DDL workloads. Moreover, to the best of our knowledge modern high speed network switches such as the ones stated before are unavailable in any public cloud for use as a test-bench. Hence, ArtISt-sim can enable researches to develop novel DDL scheduling techniques with modern networking hardware.
To summarize, in this paper, our primary contributions are outlined as follows:
  • We present Dally, a novel DL cluster scheduler, crafted to reduce end-to-end Makespan, average JCT and communication overhead. Dally employs delay scheduling to achieve optimal job consolidation and placement and implements a network-sensitive preemption policy, facilitating prioritized placements of jobs with heightened sensitivity to network conditions.
  • Dally also incorporates an ”auto-tuner” capable of dynamically adjusting the delay timers, based on network usage, required for delay scheduling.
  • Further, we develop a high-fidelity (iteration-level) DDL cluster simulator using ASTRA-sim called ArtISt-sim, which effectively emulates a DDL data center equipped with contemporary networking hardware.
  • Finally, extensive evaluation of Dally using ArtISt-sim, encompassing both batched and randomly arriving workloads of six real-world DNN models, reveals significant improvements. Dally improves makespan by over 69% compared to consolidation based policy, reduces average JCT by more than 83%, and slashes average communication overhead by 98% under congested networking conditions.
2. BACKGROUND AND RELATED WORK
2.1. Delay scheduling based on data locality
Delay scheduling is a traditional scheduling algorithm used in big-data clusters that defers the scheduling of a job to the future when a suitable/preferred job placement destination (hardware) is not immediately available. Instead of promptly accepting a resource offer from any worker machine, a job undergoes a ”delay time”, anticipating the availability of a worker node with enhanced data locality in the future. Data locality refers to the location of the input data for the job, which could be on the same machine, the same rack, or somewhere accessible over the network. The farther the input data is from the job, the longer it takes for the data to be read. System administrators set the delay time, often maintaining the default values established by cluster schedulers. Traditional big-data schedulers, like YARN (yarn, ) and Spark (spark, ), typically implement delay scheduling as a ”default” strategy for better data locality. Note, deep learning jobs typically assume the availability of input data on the local machine or through fast remote storage.
2.2. Simulation Platforms
ASTRA-sim (Accelerator Scaling for TRAining Simulator) is a comprehensive simulation platform facilitating parameterized descriptions of DNN model, system, and network fabric for end-to-end simulation of a DDL loop. ASTRA-sim users can configure various DDL training setups, including DNN models (number and types of layers, GEneral Matrix Multiplication (GEMM) size (gemm, ), etc), DDL parallelism (data, model, hybrid, etc.), fabric design (number of links, latency/bandwidth per link, etc.), and fabric topology (2D/3D torus, switch, ring, etc). ASTRA-sim leverages network-layer simulators, e.g., Garnet (garnet, ), Analytical (analytical, ), and ns3 (rashidi2020scalable, ), for packet-accurate simulations. An improved version of ASTRA-sim (won2023astrasim20, ) with capabilities such as arbitrary model parallelism, improved hierarchical topology parameterization and enhanced memory system modeling is now available. Also, (Akella22, ) developed a simulator using ASTRA-sim and ns3 to model various congestion control schemes.
ASTRA-sim expects the following three input files to simulate a single loop of DDL training:
$\bullet$ Workload file: This file contains the delay in cycles for the GEMM operations of each layer during the forward and backward pass along with layer-wise communication size.
$\bullet$ System file: The system file provides parameters to the network scheduler such as topology, packet scheduling policy (e.g., LIFO, FIFO), all reduce implementation, etc.
$\bullet$ Network file The network files encompass details regarding the physical network topology, network dimensions, link information (bandwidth, latency, etc.) etc.
We want to emphasize however that ASTRA-sim functions as a single DDL job simulator and lacks the capability to simulate a large GPU cluster running multiple DL jobs.
2.3. Prior Work on DL cluster scheduling
A plethora of DL cluster schedulers have been developed recently with varying objectives such as JCT performance, fairness, and SLO compliance. In terms of performance, approaches like (lucid-asplos23, ; ant-man, ) focus on packing multiple DNN models onto a single GPU, boosting parallelism and decreasing overall execution time. Our work, centered around network sensitivity to enhance performance, is orthogonal to such efforts, as our proposed techniques can be readily applied to GPUs simultaneously running multiple models. However, the absence of communication characterization for such ”packed models” hinders the application of our approach. Network agnostic works such as (pollux, ; optimus, ; gandiva, ) improve performance through resource reallocation, while (horus, ) is an interference-aware scheduler which is designed for a heterogeneous GPU cluster. For fairness, (themis-nsdi, ; gandiva-fair, ) have been proposed, while (chronus, ) is an SLO compliant scheduler.
3. MOTIVATION
A typical DL cluster today comprises a large number of GPUs communicating via different types of network interconnects. These interconnects may be (i) intra-machine like NVSwitch (nvlink-nvswitch, ); (ii) intra-rack connected via Infiniband; or, (iii) inter-rack connected via modern network (Ethernet) switches. Such a hierarchical network structure is depicted in Fig. 1. In this figure, a job mapped to GPUs on the same machine is shown in red; a job running on the same rack (but on different machines) is marked in blue; and a job mapped to GPUs across the network is shown in grey. The communication overhead is lowest for the jobs highlighted in red, followed by those in blue and then grey.
Figure 1. A typical (hierarchical) datacenter n/w.
3.1. Communication overhead in DDL Clusters
DDL training throughput is heavily influenced by communication speed and latency due to the significant volume of data transfer among DNN workers during gradient synchronization. As highlighted in (kungfu, ), an 8-GPU server engaged in training a ResNet50 model produces 4 GB of gradients per second. Under layer-wise model synchronization, starting from the output layers under back-propagation: (i) gradients (of the loss training-objective based on the GPUs’ assigned training-data batches) are computed with respect to parameters of the current layer, (ii) these gradients are then exchanged among the GPUs, and finally, (iii) the current layer’s parameters are simultaneously updated by each GPU. Hence, multiple DDL models can quickly overwhelm the link bandwidth. This underscores the importance of high-speed and low-latency interconnects between GPUs for mitigating communication-related stalls. Previous studies such as (pipedream, ; stash, ) reveal that the communication overheads can be as high as 5×5$\times$5 × the actual compute time. To circumvent such substantial overheads, schedulers like Tiresias strive to consolidate DDL jobs on the most proximal GPUs possible.
Figure 2. Single iteration running time for models consolidated on the same machine, rack, and across the network
3.2. Job consolidation in DL cluster
Cluster schedulers like Tiresias (tiresias, ) aim to reduce communication overhead by consolidating job placement, and assigning GPUs to a job within the fewest machines possible. However, due to multi-tenancy, such consolidation can lead to lengthy queueing delays, depending on the cluster’s load. Individual jobs may reject resource offers generated by the cluster scheduler if the offered GPUs are not located on the same machine or rack, especially when GPU demand is high. As a result, jobs remain queued, increasing their job completion time (JCT).
However, not all jobs benefit equally from consolidation, as network overhead varies across DNN models. More specifically, jobs which are less sensitive to network quality (less ‘network-sensitive’ models) can tolerate less stringent network consolidation (consolidation on physically distant GPUs) compared to highly ‘network-sensitive’ jobs. To illustrate this, in Fig. 2, we simulate the single iteration DDL running times using ASTRA-sim for various models with different network placements. As expected, in general, consolidation on physically proximal GPUs yields lower run times. However, we observe significant variations in communication overheads across network tiers for different DNN models (also shown in Table 1). Thus, model network sensitivities could act as a potentially relevant metric in determining the priority of each job with respect to consolidation: the higher the network sensitivity of a model, the more it should be prioritized to get an allocation with physically proximal (consolidated) GPUs.
In this direction, Tiresias, seeks to categorize network-sensitive jobs by employing model ”skew,” which is defined as the ratio of the largest tensor size to the overall model size. Tiresias consolidates jobs with high skew, while other jobs accept any resource offer they receive. However, using ASTRA-sim to analyze the communication overheads of various models under different network placements and high speed networking hardware reveals weak correlation between skew and network sensitivity as shown in Table 1. For e.g. ResNet18 has low skew but very high network level communication overhead and VGG11 has high skew but low communication overhead. This could be due to the rapid advancements in state-of-the-art (SOTA) network hardware.
Model Skew
Machine
Rack
Network
VGG11 (vgg, ) High 1% 6% 7%
AlexNet (alexnet, ) High 2% 13% 100%
MobileNetV3 (mobilenet, ) High 42% 940% 19592%
ResNet18 (resnet, ) Low 7% 116% 2749%
ResNet50 (resnet, ) Low 12% 12% 38%
BERT large (bert, ) Low 8% 23% 715%
Table 1. Communication overhead as a % of compute time.
3.3. Improvements in SOTA network hardware
Recent advancements in network hardware have introduced remarkably high network bandwidth with exceptionally low latency. Hardware manufacturers, such as NVIDIA, are now offering network switches that deliver several-fold improvements in QoS compared to previous-generation hardware. For intra-machine networking, NVSwitch (nvlink-nvswitch, ) provides a bandwidth of up to 900 Gb/s, surpassing PCIe’s bandwidth of only up to tens of GB/s combined with high latency (li2019evaluating, ). Furthermore, for Infiniband-based intra-rack networking, NVIDIA quantum switches now offer up to 400 Gb/s bandwidth per port with ultra-low latency using GPU direct RDMA. In contrast, standard Infiniband offers only up to 50 Gb/s bandwidth while experiencing relatively high latency (up to tens of microseconds) (Infiniband-eval, ). Finally, NVIDIA spectrum Ethernet switches provide up to 800 Gb/s bandwidth for inter-rack networking. However, current schedulers lack awareness of such high-speed network hardware, which could be harnessed for improved performance. This underscores the demand for a novel approach to DL cluster scheduling.
3.4. Limitations of current SOTA DL cluster schedulers
Current SOTA schedulers like Tiresias and Gandiva (gandiva, ) have several shortcomings such as: (i) Leveraging only limited knowledge of the data center network: due to the lack of awareness of the various network tiers, for e.g., the medium-bandwidth GPU direct RDMA that exists within a rack, these schedulers often end up making a sub-optimal job placement decision, in this case, mapping the job across multiple racks. (ii) A strict consolidation policy: restricting the job placement to either the fewest machines possible (for high skewed models) may lead to high queueing delays in case the former is not available. This necessitates an intermediate approach, allowing for a gradual and graceful ‘relaxation’ of consolidation, particularly for moderately sensitive jobs based on queueing delay and network sensitivity. (iii) Sub-optimal job priority: Tiresias employs a generalized version of Least Attained Service (LAS) (las, ) called ”Discretized 2D-LAS” (2DAS) (tiresias, ) to determine job priority. The 2DAS value for a job is calculated by multiplying the time for which it has been running by the number of GPUs it utilizes. Tiresias prioritizes jobs based on their 2DAS values, with lower values indicating higher priority. Consequently, a job with the highest 2DAS value is least likely to be scheduled for execution and most likely to be preempted out. However, a job with unfavorable network placement and high network sensitivity may have been running for an extended period but may not have made significant progress due to the excessive network overhead associated with its placement. This would lead to a high 2DAS value and make it less likely to receive resource offers that are consolidated, as jobs with lower 2DAS values would receive the most favorable offers.
These limitations of existing DL cluster schedulers underscore the need for a new approach that incorporates both network awareness and network sensitivity-based job consolidation. However, development of such scheduler for cutting-edge network technology is hindered by the high cost of the associated experimental setups and the availability of such hardware, necessitating a high-fidelity cluster simulator.
3.5. A high-fidelity simulator for DDL scheduling research
The exorbitant cost of deep learning GPUs, with several GPUs required to create a large deep-learning cluster, poses a significant challenge in conducting real-world experiments. This, combined with the need to run such experiments for extended periods, can escalate the cost to the hundreds of thousands of dollars. Even utilizing public cloud resources may prove uneconomical as deep learning machines like the AWS P4 instances incur hourly charges of more than \$32/hour (aws_pricing, ). Further complicating matters, research based on cutting-edge network hardware necessitates a simulation platform due to its limited availability. To the best of our knowledge, no public cloud provider offers access to such hardware. To address these challenges, research efforts like (rashidi2022themis, ) have turned to ASTRA-sim as a viable tool for simulating the latest network hardware. However, ASTRA-sim is a single DDL job simulator.
Existing multi-job DDL cluster scheduler simulators, such as (themis-nsdi, ), suffer from low accuracy due to their unrealistic approach to network overhead modeling. For example, (themis-nsdi, ) implements static slowdown penalties of job placement on the rack and the network irrespective of the DNN model or network hardware. But to achieve high confidence in the simulation results, a DL cluster scheduler simulator must dynamically determine the accurate network overhead of a job based on specific placement within the cluster and the type of network hardware used to accurately reflect real-world conditions. In response to this requirement, we propose a multi-job DL cluster simulator that employs ASTRA-sim to accurately determine network overhead.
4. DALLY DESIGN
We now describe Dally, a novel network-placement sensitive DL cluster scheduler that leverages the principles of delay scheduling and novel network-sensitive preemption policy. In the following sections, we delve into the details of our system design, including the simulation methodology employed to evaluate its effectiveness.
4.1. Overview
4.1.1. System goal
The system processes a stream of DL/DDL jobs that need to be scheduled for execution. These jobs exhibit heterogeneity in terms of the DNN model, the number of iterations required to attain the target accuracy, the compute time per iteration, and the GPU demand per job. The system may encounter jobs arriving in batches so that the aggregate GPU demand across all jobs would normally surpass the total number of GPUs available in the cluster. Alternatively, jobs may arrive according to a Poisson distribution. Our goal is to construct a DL cluster scheduler that aligns with the target objectives and assumptions elaborated upon in the subsequent sections. This scheduler will function within a cluster comprising homogeneous GPUs connected via latest-generation network hardware in a hierarchical structure, as discussed in Section 3.
4.1.2. Objective
The overarching goal of our scheduler is to minimize cloud expenditure for its users by reducing the time spent occupying cluster GPUs. This objective is achieved through two primary strategies.
  1. (1)
    Minimizing the end-to-end Makespan for training multiple DNN models when submitted as a batch.
  2. (2)
    Reducing job completion time (JCT) for individual jobs.
4.1.3. Constraints
Dally must achieve the aforementioned objectives subject to the following constraints:
  1. (1)
    Unknown JCT distributions: We assume that the prior running time of any job is unknown.
  2. (2)
    Absence of model architecture information: The scheduler lacks the knowledge of the specific DNN model architecture associated with each job.
  3. (3)
    Unknown job submission pattern: The scheduler lacks information about the exact order and frequency of job arrivals.
  4. (4)
    Unknown GPU demand: The GPU resource requirements for individual jobs and the overall GPU demand across all running jobs are unknown beforehand.
4.1.4. System workloads
We employ the cluster job trace made available by SenseTime (sensetime, ) through the work presented in (lucid-asplos23, ). The traces are more recent than those utilized in (philly_trace_analysis, ) and provide a more realistic representation of a real-world DL cluster with multiple jobs being submitted over an extended period. This trace encompasses individual job arrivals alongside pertinent metadata information for each job.
4.2. Simulation methodology
In this work, we develop a high-fidelity DDL cluster simulator – Artificial Intelligence System Simulator (ArtISt-sim) – capable of accurately replicating network slowdowns induced by specific job placements throughout a job’s execution.
4.2.1. Simulator design
ArtISt-sim builds upon a prior work, Themis (themis-nsdi, ), which is a DL cluster scheduler simulator, and ASTRA-sim. The choice of Themis is based on its unique capability to simulate DL jobs at an iteration level, i.e., a feature not present in other simulators like (lucid-asplos23, ; chronus-socc21, ; tiresias, ). This granular simulation approach enables us to precisely track ”job progression” for preemption and subsequent placement decisions throughout the job’s lifetime. However, as previously discussed, the network slowdown in the original work is encoded statically. To address this limitation, in ArtISt-sim, we modify Themis to invoke ASTRA-sim for each new job placement (refer Fig. 3).
For executing ASTRA-sim with specific network placements, ArtISt-sim dynamically generates the necessary input – workload, system, and network topology files and passes them to ASTRA-sim (blue arrow in Fig. 3). These generated files delineate the ”portion” of the data-center network utilized by the worker GPUs. For instance, in Fig. 1, the grey colored job occupies 8 GPUs across two machines connected via a network link. To simulate ASTRA-sim for such a placement, ArtISt-sim produces the 8-GPU configuration files with a two-tier hierarchy representing the placement. The configuration files specify that the intra-machine GPUs are connected via NVSwitch in a ring topology, and the two machines are linked via a network (Ethernet) switch using a switch topology. Other ASTRA-sim configurations, such as collectives, are also dynamically generated. ASTRA-sim then accurately simulates the job for a single DDL iteration based on the placement and relays the communication overhead back to ArtISt-sim (blue arrow in Fig. 3), which subsequently utilizes this communication overhead to simulate the job’s progression (iterations) until it is either interrupted due to preemption or successfully completes its execution. Note that the compute time per iteration is extracted from the workload trace, and ASTRA-sim serves solely to calculate the network overhead for the specific placement. Network overhead is calculated by first finding the percentage of exposed communication time to the compute time in the ASTRA-sim simulation. The percentage value is applied to the iteration time from the trace to calculate the network overhead and the actual iteration time in ArtISt-sim.
Figure 3. Simulation design
4.2.2. Simulation calibration and accuracy
To ensure the fidelity of our simulations, we employ a calibration process that incorporates real-world measurements into ArtISt-sim. We execute each model listed in Table 1 on an 8-GPU machine connected via NVSwitch and calculate the associated network overhead. Along with this, we simulate the model with the same network configurations (NVSwitch) in ASTRA-sim. Subsequently, we scale the workload configuration file of ASTRA-sim (for each model) to eliminate the discrepancy between simulated and actual network overhead. These modified workload files are then used for all ASTRA-sim calls. The ”calibrated” simulator is now ready to simulate overhead arising from rack and network-level placements.
For the verification of the various baseline scheduler implementations in ArtISt-sim, we rely on the fact that Themis has already been validated in (themis-nsdi, ). Furthermore, we faithfully reproduce the design principles of the Tiresias and Gandiva schedulers as outlined in their respective publications.
4.2.3. Implementation
We have implemented ArtISt-sim in about 2000 lines of java. [^2] Upon acceptance of this work, we will freely provide comprehensive details of all configurations along with the open-source code. ArtISt-sim requires the network configuration parameters for intra-machine, intra-rack, and inter-rack interconnects. These parameters encompass the network topology, bandwidth, latency, collective communication algorithm, and other relevant details. Additionally, the user needs to specify the number of machines per rack and the total number of racks in the cluster.
4.3. Design of the Dally scheduler
Dally is designed to minimize cloud expenditures in both public and private cloud environments. To achieve this objective, Dally uses delay scheduling for job placement and employs a novel preemption strategy that incorporates network sensitivity. When a job is submitted to the cluster, the global scheduler (refer to Fig. 4) initially places it into a wait queue. The scheduler then periodically generates ”resource offers” to jobs in the wait queue based on the availability of cluster resources. If a job’s local scheduler (Fig. 4) decides to accept the offered resources, it is moved to the run queue, where it executes until it is either preempted or completes its execution. Upon preemption, the job is reintroduced into the wait queue and remains there until it accepts another resource offer. An instance of Dally’s mechanisms governing preemption and job placement is elaborated upon in the following sections.
Figure 4. Scheduling scheme.
4.3.1. Preemption priority
As already discussed in Section 3, Tiresias’s 2DAS based priority mechanism is oblivious to network slowdowns. In a large networked cluster, an effective priority mechanism should consider the relative network sensitivity of jobs and the slowdowns they have experienced. Therefore, we propose a network-sensitive priority metric to address this shortcoming. We define the number of iterations completed as $I_{compl}$ and the total expected iterations as $I_{total-expected}$. Our approach necessitates awareness or prediction of the overall expected iteration count. This information can be readily obtained from openly accessible benchmarks like MLperf (mlperf_mlsys20, ). Alternatively, users can estimate the total iteration count needed for convergence through established extrapolation techniques such as (optimus, ). It is important to emphasize that only an approximate value for the total expected iterations suffices.
Next, we define $T_{run}$ as the time spent by the job in the run queue and the total ideal job running time $T_{total-ideal-run}$ as the compute time per iteration of a job multiplied by the total number of expected iterations. The compute time per iteration can be measured by running a single training iteration of the model on a single GPU, where the iteration comprises the forward pass and backward pass (we assume no communication overhead for ideal JCT calculations). We now define work completed $W_{compl}$ and normalized running time $T_{norm}$ as:
$W_{compl}=\frac{I_{compl}}{I_{total-expected}}\qquad T_{norm}=\frac{T_{run}}{T_{total-ideal-run}}.$
Finally, we define our network sensitive priority $Nw_{sens}$ as:
$Nw_{sens}$ $=\frac{W_{compl}}{T_{norm}}.$
A lower value of $Nw_{sens}$ indicates that the job has endured greater network-induced slowdowns, and, consequently, its priority should be increased. Resource offers are hence made in increasing order of $Nw_{sens}$. This allows jobs with lower $Nw_{sens}$ to have a higher probability of receiving more consolidated resource offers.
4.3.2. Job placement
In a DL cluster, jobs may receive resource offers with varying levels of consolidation, meaning the GPUs offered may be located on a single machine, within a single rack, or spread across the network. Dally employs the well-established delay scheduling strategy to handle these offers. Under this scheme, a job has the option to decline offers that do not meet its consolidation preference as described in Algo. 1. Initially, a job may reject all offers that do not place its GPUs on the same machine until a machine-level delay timer has elapsed while waiting (starving) for resources (line 10 in Algo. 1). Once the machine-level delay timer elapses, the job will commence accepting offers that position all its GPUs within a single rack until the rack-level delay timer has also elapsed (line 18 in Algo. 1). Subsequently, at this point, the job will accept any resource offer without considering consolidation (line 21 in Algo. 1). Note that a job that cannot fit on a single machine will have a machine-level delay timer set to 0, and similarly, a job that cannot fit on a single rack will have both delay timers set to 0.
The underlying principle is to delay the placement of jobs in anticipation of better consolidation options becoming available in the future. This delay-based approach strikes a balance between the high queuing times associated with consolidation and the high communication overhead of non-consolidation. In traditional big data schedulers, these delay timers are typically configured by system administrators and often left at their default values of a few seconds. Based on our empirical observations, we establish a default value of 12 hours for machine-level consolidation and another 12 hours (or 24 hours in total) for rack-level consolidation. Delay timers need to be adjusted based on resource contention in the cluster for optimal delay scheduling. Generally, system administrators adjust timers based on historical contention data of the cluster. But, to eliminate the need for manual timer adjustments, we introduce an ”auto-tuner” that automatically optimizes these delay timers. We elaborate on our auto-tuner in the following section.
4.3.3. Auto-tuner
We introduce an auto-tuner module in Dally aimed at automatically fine-tuning delay timers. This tuner relies on ”moving averages” derived from historical job waiting (starvation) times associated with specific consolidation and GPU demands. The core concept revolves around leveraging historical data to determine an optimal waiting (delay) duration for securing a favorable job consolidation. When a job accepts a resource request, it updates a ”global list” pertaining to the consolidation level (machine or rack) and the job’s GPU demand, with the time it waited for before accepting a resource offer (lines 7 and 15 in Algo. 1). Dally maintains comprehensive lists, tracking job waiting times for each combination of consolidation and GPU demand. In a DL cluster, the GPU demands usually align with multiples of 2, typically ranging from 5 to 10 types. As a result, this approach leads to a manageable 2-20 lists for machine-level and rack-level consolidation paired with GPU demands.
To calculate the moving averages (described in Algo. 2), Dally mandates users to define the size of the wait-time list, denoted as the ”history limit” (sliding window size) for the auto-tuner. The maintenance of the wait-time lists involves discarding values (utilized in average calculation) that exceed the specified history time limit (lines 4 and 10 in Algo. 2). Whenever a job receives a resource offer, the delay timer is calculated by averaging the wait-time list corresponding to the consolidation level (machine or rack) and GPU demand of the job, and then adding two sample standard deviations (similarly estimated using the same wait-time list). This value is then used in determining the optimal duration for a job to wait for a favorable consolidation (lines 13 and 14 in Algo. 2). We present the timeline depicting the adjustment of delay times (rack level) for a typical ¿P95 DDL job in Fig. 5. It is evident from our observations that the auto-tuner modifies delay timers by increasing the value when there is a lower number of racks (and consequently fewer GPUs) in the cluster leading to more resource contention. Over time, the auto-tuner effectively ”learns” the optimal duration to wait as demand delays are updated.
Figure 5. Auto tuning timeline
Note that it is essential to tailor the history limit according to the cluster size. Specifically, in larger clusters, a smaller history time limit is advised. This is simply because more jobs get placed over time in a large cluster compared to a smaller cluster.
Algorithm 1 On Resource Offer
1:Resource offer $R$
2:Accept or reject offer
[…]
4:$T_{starvation}=T_{curr\_time}-T_{last\_assigment}$
5:$g_{d}=\mbox{GPUs demanded by the job}$
[…]
7:if $g_{d}\leq\mbox{GPUs available on a single machine in }R$ then
8:     Allocate $g_{d}$ GPUs on machine
[…]
10:     return accept resource
11:end if
12:if $T_{starvation}<T_{Mc}$ then
13:     return reject offer
14:end if
15:if $g_{d}\leq\mbox{GPUs available on a single rack in }R$ then
16:     Allocate $g_{d}$ GPUs on rack
[…]
18:     return accept resource
19:end if
20:if $T_{starvation}<T_{Rk}$ then
21:     return reject offer
22:end if
23:Allocate $g_{d}$ GPUs on network
24:return accept offer
Algorithm 2 Get Tuned Timers
1:GPU demand $G_{d}$
2:Tuned delay timers
[…]
4:for time in mc_starvation_times do
[…]
6:         remove time from mc_starvation_times
7:     end if
8:end for
[…]
10:for time in rack_starvation_times do
[…]
12:         remove time from rack_starvation_times
13:     end if
14:end for
[…]
5. EVALUATION
In this section, we assess Dally through extensive simulations driven by large-scale traces, following the previously outlined methodology. It is important to emphasize that our evaluations leverage the most recent generation of network hardware, necessitating the use of simulations as the exclusive option.
5.1. Experimental setup
5.1.1. Workloads
In our experiment, we employ real-world production traces obtained from SenseTime. We randomly select 500 DL/DDL jobs for batch and about 400 DL/DDL jobs for Poisson workloads from the trace, [^3] From the original SenseTime trace, we replace EfficientNet with AlexNet for a high skew workload which utilize the models specified in Table 1. The table encompasses a diverse range of network overheads, providing ample coverage for testing a DDL cluster executing various network-sensitive DNN models. The cluster trace furnishes the compute requirements for individual jobs, including the time per iteration, the necessary number of iterations, and the quantity of GPUs demanded. We assume a data-parallel distributed training paradigm.
5.1.2. Cluster setup
In our experiments, we adopt the configuration of a three-tier network data center. Each machine is equipped with eight GPUs interconnected through NVIDIA NVSwitch. Within each rack, eight machines are connected via the NVIDIA Quantum switch, and the racks across the cluster are interlinked through the NVIDIA Spectrum network switch. To set up our network simulations in ASTRA-sim, we rely on the QoS data advertised by NVIDIA (nvidia-quantum, ; nvidia-spectrum, ). Our experimentation involves manipulating the number of racks, specifically using 2, 4, 8, and 16 racks, to observe the impact of varying compute and network resources.
Figure 6. Batch arrival results: (a) Makespan; (b) Tail queueing delay; (c) Average communication overhead
5.1.3. Baselines
We use the following three baselines to evaluate our scheme:
  1. (1)
    Gandiva: Being network-agnostic, Gandiva, as one of the initial DL cluster schedulers, exhibits sub-optimal performance. It addresses this limitation by migrating jobs to more favorable GPU consolidation whenever resources become available.
  2. (2)
    Tiresias: The state-of-the-art DL cluster scheduler currently employed adopts a strategy of consolidating job placement based on their concept of skew in DNN models. This scheduler possesses partial knowledge of the network, and the consolidation is stringent for skewed DNN models.
  3. (3)
    Dally-manual: This represents a modification of our proposed scheme where delay timers for delay scheduling are manually configured. This aligns with the approach utilized by contemporary state-of-the-art big-data cluster schedulers, such as YARN. Note that the preemption mechanism remains consistent with our proposed scheme Dally with auto-tuning. We use this baseline to demonstrate the effectiveness of our delay timer auto-tuner.
5.1.4. Performance metrics
We define the following metrics to measure performance benefits.
  1. (1)
    Makespan: The time elapsed between arrival of the first job to the completion of the final job in the cluster.
  2. (2)
    JCT: The job completion time (JCT) is the elapsed time between job arrival and completion in the cluster.
  3. (3)
    Queueing delay: Time spent by a job in the wait queue.
  4. (4)
    Communication overhead: Exposed communication time of a job.
5.2. Results
(a) 2 racks
(b) 4 racks
(c) 8 racks
(d) 16 racks
Figure 7. JCT CDF - Batch arrival
(a) 2 racks
(b) 4 racks
(c) 8 racks
(d) 16 racks
Figure 8. JCT histogram (P95) - Poisson arrival
5.2.1. Makespan
Fig. 6 (a) illustrates the improvement in Makespan. Notably, Dally exhibits an improvement of up to 69% compared to Tiresias and up to 92% compared to Gandiva. This improvement is attributed to the implementation of delay scheduling, effectively relaxing stringent consolidation, thereby decreasing queueing delay and optimizing JCT by judiciously awaiting resource offers with favorable consolidation. A loosening essentially implies a job being positioned in less favorable network tiers, a scenario which may not be feasible in Tiresias. The manual configuration of delay timers in Dally-manual results in sub-optimal waiting, as it cannot dynamically ”tune” its delay timers based on contention. The static nature of the delay time fails to consider GPU demand and resource contention in the cluster, leading to inferior performance compared to Dally, while still outperforming Tiresias and Gandiva due to reduced queuing delays.
5.2.2. Queueing delay
The primary contributor to the significant improvement in Makespan is the reduction in queueing delays experienced within the cluster. Although Dally reports an average queueing delay improvement ranging from 18% to 31%, depending on the cluster size, the cluster’s Makespan is more notably influenced by the ¿P95 and ¿P99 JCTs (i.e. the 95th and 99th complementary percentiles of JCT distribution). Dally demonstrates a remarkable improvement of 58% – 71% in ¿P95 and 67% – 78% in ¿P99 queueing delay compared to Tiresias (Fig. 6 (b)), which results in the substantial improvement in Makespan. The reduction in queueing delay can be attributed to the less stringent consolidation policy of Dally compared to Tiresias, although not as relaxed as Gandiva’s policy. Yet, Dally improves upon Gandiva in tail (¿P95 and ¿P99) queueing delay by as much as 81%. In fact, Gandiva’s network-agnostic scheduling policy leads to very high communication overhead.
5.2.3. Communication overhead
Dally demonstrates the lowest communication overhead, surpassing the nearest network-aware baseline, Tiresias, by a notable margin ranging from 53% to 83% (Fig. 6 (c)). This is due to Dally’s preemption policy, which prioritizes providing better-consolidated placements to jobs suffering from sub-optimal placements or network sensitivity, thereby mitigating communication overhead. Gandiva exhibits a slight performance difference in Makespan compared to Dally in the initial 2-rack experiment but experiences a substantial degradation as the number of racks increases. This discrepancy arises because the 2-rack experiment involves fewer network connections and consequently fewer network placements. However, with an increase in the number of racks, the cluster experiences a rise in network connections and leads to jobs being placed on the network, resulting in elevated communication overhead, as depicted in Fig. 6 (c). Additionally, this phenomenon results in a long tail in the number of job completions within the cluster, subsequently leading to low cluster (GPU) utilization as depicted in Fig. 9.
(a) Batch
(b) Poisson
Figure 9. Cluster GPU utilization (8 racks)
(a) Batch
(b) Poisson
Figure 10. Number of jobs remaining (8 racks)
5.2.4. Cluster utilization
We explain this section through results from the 8 racks experiment but we get similar results for the other experiments as well. Fig. 10 illustrates the number of remaining jobs in the cluster, highlighting Gandiva’s notable tailing effect. Gandiva jobs, (including the ¿P95 and ¿P99) experience substantial communication overhead, resulting in diminished cluster utilization, as depicted in Fig. 9. Dally’s efficacy in minimizing communication overhead and queuing delays contributes to the highest job completion rates among all tested schemes as shown in Fig. 10 (a proxy for cluster utilization), along with the lowest JCTs.
5.2.5. Job completion time
To assess the JCT, we utilize traces with both batch and Poisson job arrivals, recognizing that the average JCT can vary significantly based on the job arrival pattern. In Fig. 11(a), we present the average JCTs with batch arrivals, revealing that Dally achieves an improvement in average JCT ranging from 19% to 36% compared to Tiresias and 23% to 51% compared to Gandiva. Examination of the JCT Cumulative Distribution Functions (CDFs) in Fig. 7 indicates that Dally reduces JCT for a majority of the job sizes. Notably, in the 2 racks experiment, Dally experiences a 39% degradation against Gandiva, which performs better with fewer aggregate network placements. However, as the number of racks (and network connections) increases, Gandiva’s performance degrades rapidly due to increased communication overhead. Additional statistics for the JCT in the 8-rack experiment are presented in Table 2 showcasing Dally’s enhancements in tail JCT.
To assess Dally under Poisson arrival conditions, we conduct experiments analogous to those performed with batch arrivals, but with jobs exhibiting typical Poisson arrival times. Dally demonstrates improvements over Tiresias ranging from 16% to 34% and over Gandiva by 23% to 51% as shown in Fig. 11(b). However, in the 4-rack scenario, Dally experiences degradation compared to Tiresias due to the random arrival pattern inherent in a Poisson process. A drawback worth noting is that Tiresias exhibits superior performance in identifying optimal placements for this particular arrival pattern and network size (racks), improving upon its own average JCT in the 8-rack and 16-rack experiments. This indicates that the specific arrival pattern was well suited for Tiresias in the 4-racks case. Also, recall from Sec. 4.3.1 that Dally does not emphasize JCT when prioritizing jobs for GPU consolidation. However, Fig. 8 presents histograms of JCTs (P95) with Poisson arrival times, illustrating Dally’s improvement in the JCTs of the majority of the jobs. Detailed JCT statistics for the 8-rack Poisson experiment are provided in Table 3. We find Dally outperforming Tiresias in both average and median values by 35% and 38% respectively while being within 10% of the tail. Dally also outperforms Gandiva in average and median JCT by 83% and 88% respectively and tail JCT by 76-92%.
(a) Batch arrival
(b) Poisson arrival
Figure 11. Average JCT
Gandiva Tiresias Dally-manual Dally
Average 5828719 4461998 3128566 2828955
Median 183717 2547080 2552189 2401390
¿P95 17485823 17841606 7097646 5983042
¿P99 39214710 25696135 8868401 8395214
Table 2. JCT batch statistics in seconds (8 racks)
Gandiva Tiresias Dally-manual Dally
Average 17371514 4329941 4960727 2831880
Median 22722917 4212795 2787855 2595529
¿P95 71362014 15059080 39298828 16469069
¿P99 208939883 15390860 39817762 16681241
Table 3. JCT Poisson statistics in seconds (8 racks)
6. CONCLUSIONS
Training of DDL jobs on GPU clusters is expensive either in a public or private cloud environment. A significant part of this time is due to the communication overhead between a hierarchical organization of GPUs in a datacenter. This communication overhead and the resultant overall training time can be minimized by a prudent allocation of GPUs, where a distributed training job is consolidated on physically proximal GPUs in proportion to its network sensitivity.
In this context, we propose Dally, a cost-effective DL cluster scheduling scheme which employs delay scheduling for job placement and consolidation. Dally incorporates a novel network-sensitive preemption technique to prioritize jobs with high network sensitivity, ensuring they receive more consolidated job offers. Additionally, we propose and integrate an auto-tuner into Dally to optimize delay timers according to varying GPU contention/demand for delay scheduling. To assess the performance of Dally, we introduce ArtISt-sim, a cost-effective DL simulation platform tailored for simulating large DL clusters. Through extensive trace-driven simulations, we demonstrate that Dally can reduce end-to-end makespan time by 69% compared to consolidation based policy, average JCT by up to 83%, and communication overhead by up to 98% under congested networking conditions.
Acknowledgements
This research was supported in part by NSF grants #2116962, #2122155, #2028929.